How many grams of 24k gold do I need to melt with 10 grams of 10k gold to make 14k gold?

Answer by checkley79(3341)

24x+10*10=14(10+x)
24x+100=140+14x
24x-14x=140-100
10x=40
x=40/10
x=4 grams of 24 caret gold is needed.
Proof:
24*4+10*10=14(10+4)
96+100=14*14
196=196